# JMI at SemEval 2024 Task 3: Two-step approach for multimodal ECAC using in-context learning with GPT and instruction-tuned Llama models

### v1

@@ 1 Introduction
Emotion Cause Analysis (ECA) is centered around the extraction of potential cause clauses or pairs of emotion clauses and cause clauses from human communication, enabling a deeper understanding of communication dynamics. By incorporating multimodal cues like visual scenes, facial expressions, and vocal intonation, it facilitates a comprehensive and technically robust analysis of the factors that trigger diverse emotional reactions (Mittal et al., 2021; Zhang and Li, 2023; Zheng et al., 2023b). Despite the considerable amount of research conducted using diverse audio, visual, and text modalities (Gui et al., 2018; Xia and Ding, 2019; Fan et al., 2020; Shoumy et al., 2020; Abdullah et al., 2021), there has been a noticeable gap in the exploration of multimodal ECA in natural settings (human conversations). In this context, Wang et al. (2023a) introduce Multimodal Emotion Cause Analysis in Conversations (ECAC) task and provide Emotion-Cause-in-Friends (ECF) dataset, which incorporates text, audio, and video modalities. This task consists of two sub-tasks: Textual Emotion-Cause Pair Extraction in Conversations (Subtask 1) and Multimodal Emotion Cause Analysis in Conversations (Subtask 2). A detailed description of these sub-tasks can be found in the task description paper (Wang et al., 2023a).
In our submission to Subtask 2 of multimodal ECAC, this paper presents two distinct approaches to address the ECAC problem, giving competitive results. Drawing inspiration from the effectiveness of LLMs in diverse downstream tasks (Wang et al., 2023b, 2024; Yang et al., 2024), including emotion recognition, we propose two LLM-based approaches that decompose the emotion-cause pair extraction process into two steps. The first step involves predicting the emotions of the utterances in the conversation. In the next step, we utilize these emotion labels to guide cause extraction. Approach 1 involves instruction-tuning two separate Llama 2 models for emotion and cause prediction, while Approach 2 leverages the in-context learning (ICL) capabilities (Dong et al., 2023) of the GPT-3.5 model. Additionally, we introduce an efficient technique using the GPT-4V model to extract conversation-level descriptions from video modality.
@@ 2 Background
Dataset
@@
We use the Emotion-Cause-in-Friends (ECF) dataset provided by Wang et al. (2023a), which is summarized in Table 3 in Appendix B. This dataset contains 13,509 multimodal utterances that occur in the American sitcom Friends with 9272 emotion-cause pairs. Each utterance consists of the text, video, and audio.
3 Methodology
3.1 Overview
We treat the task at hand as a two-step process. In the first step, we predict the emotion of each utterance in all N conversations. Here, the context $C_{j}$ for utterance $u_{j}$ of conversation $D_{i}$ is the entire conversation itself. Given $E$ target emotion labels and $\hat{y_{j}^{e}}$ as the predicted emotion label, the problem can be formulated as:
$\hat{y_{j}^{e}}=\arg\max_{e}\mathcal{P}(y^{e}|u_{j},C_{j},\theta)$ (1)
where $\theta$ denotes the parameters.
In the second step, given these emotion labels, we predict the causes of each utterance that has an emotion other than neutral. The causes will be a subset of all utterances in the conversation $D_{i}$. Let the learned function be $f:U\rightarrow 2^{U}$, where $U$ is the set of all utterances in the given conversation. It predicts the subset $\hat{y_{j}^{c}}$ of cause of emotion utterance $u_{j}$ as:
where $\hat{y_{j}^{e}}\neq$ neutral.
3.2 Approach 1: Fine-tuned Llama-2
In our first approach, we perform instruction fine-tuning of the Llama 2 Large Language Model, an open-source model developed by GenAI, Meta (Touvron et al., 2023). From the three variants with 7, 13, and 70 billion parameters, we use the 13 billion parameter model due to resource constraints, albeit the performance of this model achieves state-of-the-art results on various downstream NLP tasks compared to other models of similar sizes. In addition, we use the Llama 2-chat version of the model [^1] https://huggingface.co/meta-llama/Llama-2-13b-chat-hf, which is optimized for dialogue use cases as it aligns with our task. In our approach, we use Llama2 API [^2] https://www.llama2.ai/ for prompt engineering. Through zero-shot prompting, we select optimal prompts for emotion identification and cause prediction. We observed that treating these two tasks separately resulted in better model output. This approach involves first identifying the emotions of all utterances in the conversation. We then add these emotion labels to the conversation and prompt the model to predict the causes for each emotion utterance. Consequently, we perform supervised fine-tuning of two separate Llama 2 models for these tasks. Although this increases the inference time, the significant performance gains outweigh the introduced latency. We treat both tasks as conditional generation, where the model generated the emotion label in the first case and the cause list in the second case, given the prompt. Detailed explanations of these approaches are provided in the following sections. The fine-tuning procedure is shown in Fig.1.
Figure 1: Pipeline for fine-tuning Llama (Approach 1)
3.2.1 Emotion recognition
To perform emotion recognition, we create a dataset where each sample includes an utterance $u_{j}$ from one of the $N$ conversations $D$ for which the LLM needs to output the emotion label. We incorporate the entire conversation $D_{i}$ along with speaker information as context in our prompt. This contextual information enhances the model’s understanding of the flow of emotions within the conversation, as demonstrated by our ablation studies in Section 5. The instruction $I_{j}^{e}$, which gave the best results, is given in Appendix  F.1 along with detailed prompt examples. The prompt consists of the instruction $I_{j}^{e}$ and the context $C_{j}$ for utterance $u_{j}$:
$Prompt_{j}=(C_{j},I_{j}^{e})$ (3)
$\hat{y_{j}^{e}}=\textbf{Llama}_{e}(Prompt_{j},\theta)$ (4)
3.2.2 Cause prediction
To prepare the dataset for cause prediction, we incorporate the emotion labels obtained for each utterance. The conversation contect now includes the emotion labels for each utterance $u_{j}$ excluding those with a predicted emotion label $\hat{y_{j}^{e}}$ of neutral. This approach enhances the model’s ability to analyze causal dependencies and identify which utterances may have contributed to a specific emotion. The output for cause prediction is a list of cause utterance IDs. The instruction is provided in Appendix  F.1. The modified prompt for this step consists of this instruction $I_{j}^{c}$ along with the conversational context containing emotion labels $C_{j}^{e}$:
$Prompt_{j}=(C_{j}^{e},I_{j}^{c})$ (5)
$\hat{y_{j}^{c}}=\textbf{Llama}_{c}(Prompt_{j},\theta)$ (6)
3.2.3 Adding video captions
3.3 Approach 2: In-Context-Learning GPT
Figure 2: Pipeline of In-Context-Learning GPT Method (Approach 2)
3.3.1 Video Captioning
Figure 3: Video Captioning Pipeline
3.3.2 Emotion Recognition
GPT tends to be uncontrollable when performing zero-shot recognition of emotions in conversations Qin et al. (2023) outputting emotions that are not a valid category of labels. To guide and control the process, we leverage in-context learning (ICL) by retrieving a conversation from the training set whose emotions are already annotated. The emotions in this conversations are explained by GPT-3.5 (Appendix Fig. 17). This retrieved conversation and its explanation serve as a demonstration for GPT to learn from, enabling it to recognize emotions in conversations more accurately. In addition, the prompt template includes the video caption as part of the input, as shown in Appendix Fig. 18.
To ensure effective ICL, it is important to provide general and descriptive examples that aid in solving the current task. In our approach, we sampled conversations from the training set containing all emotion categories. These conversations were stored as text-embedding-ada-002 embeddings (Neelakantan et al., 2022) in a vector database. At test time, we compute the embedding for a conversation and retrieve the closest matching embedding from the database based on Euclidean distance. The retrieved embedding aids ICL in improving emotion understanding and recognition.
3.3.3 Cause Prediction
Following prediction of emotions, we proceed to predict the causes for each emotional utterance within a context window around that utterance. The bounds of the context window are discussed in Appendix C.2 We chose not to predict causes across entire conversations to save costs on tokens.
For predicting causes of an utterance with emotion $e$ within a given context window $c$, we retrieve context windows containing utterances with the same emotion $e$ that exhibit semantic similarity to $c$. This retrieval is accomplished through the Euclidean distance comparison of text-embedding-ada-002 embeddings derived from the training data. The retrieved conversation’s causes are explained by GPT-3.5 (Appendix Fig. 19). Learning from the explained retrieved-context windows, cause prediction on $c$ can be performed by GPT-3.5. Video captions are also included in the prompt (Appendix Fig. 20), since the local window may have lost some broader context.
4 Experimental setup
Training details
For approach 1, the data is split into train, test, and validation sets in the ratio 8:1:1. We use peft library  [^4] https://huggingface.co/docs/peft/en/index for Parameter-Efficient Fine-Tuning. Due to memory constraints, we fine-tune a 4-bit quantized Llama-2 model using bitsandbytes library  [^5] https://github.com/TimDettmers/bitsandbytes. We report the details of the implementation for both approaches in Appendix  C.
Evaluation metrics
5 Results and Discussion
Main results
System w-avg F1 F1
1. Samsung Research China-Beijing 0.3774 0.3870
2. NUS-Emo 0.3460 0.3517
3. SZTU-MIPS 0.3435 0.3434
4. GPT-ICL (Ours) 0.2758 0.2816
5. MotoMoto 0.2584 0.2595
6. Fine-tuned Llama (Ours) 0.2558 0.2630
Table 1: Leaderboard Results on Evaluation Data
Ablation study
We conduct extensive ablation studies to measure the importance of the techniques we employ summarized in Table  2. For these experiments, we use a subset of our test set from the training set containing 528 utterances. It can be seen that the performance of zero-shot Llama as well as GPT is the lowest. Instruction-tuning and ICL clearly improve the performance on the task, showcasing the significance of making LLMs context-aware when tackling downstream tasks. Adding self-causes improves performance in both zero-shot and context-aware cases highlighting their importance. The detailed table is in Appendix  D.
Approach F1 w-avg F1
Zero-shot Llama
    - w/o self-causes 0.117 0.116
    - w/ self-causes 0.222 0.215
Instruction-tuned Llama
    - w/o self-causes 0.325 0.318
    - w/ self-causes 0.364 0.352
Zero-shot GPT
    - w/o self-causes 0.100 0.097
    - w/ self-causes 0.189 0.184
In-context-learning GPT
    - w/o self-causes 0.235 0.241
    -w/ self-causes 0.329 0.334
Table 2: Results on Validation Set.
Limitations
6 Conclusion
Appendix A Related Work
Text-based ECA
Multimodal ECA
Large Language Models
Appendix B Dataset
We use the Emotion-Cause-in-Friends (ECF) dataset curated by Wang et al. (2023a) summarized in Table 3. This dataset contains a total of 13,509 multimodal utterances that occur in the American sitcom Friends with 9272 labeled emotion-cause pairs. Each utterance consists of the text with corresponding video and audio.
@@ Appendix B Dataset
Appendix C Implementation details
C.1 Training details for Llama
@@
C.2 Details for in-context learning GPT
@@
Context windows
The bounds of the context window are given in Table 6. The bounds were informed by the distribution of the majority of relative positions of causes in the training set Appendix B).
Position Previous Next
Beginning 0 2
End 5 0
Middle 5 2
Table 6: Context Window Bounds in each Direction
@@
C.3 Post-Processing
In both our approaches, after getting the causes, we perform a post-processing step where we add the emotional utterance as its own cause which we call self-causes. This gives significant performance boosts as a majority of the causes are self-causes as pointed out in Appendix  B.
Appendix D Detailed results
@@
Appendix E Error Analysis
@@
Appendix F Prompt details
F.1 Fine-tuned Llama 2

### v2

@@ 1 Introduction
Emotion Cause Analysis (ECA) is centered around the extraction of potential cause clauses or pairs of emotion clauses and cause clauses from human communication, enabling a deeper understanding of communication dynamics. By incorporating multimodal cues like visual scenes, facial expressions, and vocal intonation, it facilitates a comprehensive and technically robust analysis of the factors that trigger diverse emotional reactions (Mittal et al., 2021; Zhang and Li, 2023; Zheng et al., 2023b). Despite the considerable amount of research conducted using diverse audio, visual, and text modalities (Gui et al., 2018; Xia and Ding, 2019; Fan et al., 2020; Shoumy et al., 2020; Abdullah et al., 2021), there has been a noticeable gap in the exploration of multimodal ECA in natural settings (human conversations). In this context, Wang et al. (2023a) introduce Multimodal Emotion Cause Analysis in Conversations (ECAC) task and provide Emotion-Cause-in-Friends (ECF) dataset, which incorporates text, audio, and video modalities. This task consists of two sub-tasks: Textual Emotion-Cause Pair Extraction in Conversations (Subtask 1) and Multimodal Emotion Cause Analysis in Conversations (Subtask 2). A detailed description of these sub-tasks can be found in the task description paper (Wang et al., 2024a).
In our submission to Subtask 2 of multimodal ECAC, this paper presents two distinct approaches to address the ECAC problem, giving competitive results. Drawing inspiration from the effectiveness of LLMs in diverse downstream tasks (Wang et al., 2023b, 2024b; Yang et al., 2024), including emotion recognition, we propose two LLM-based approaches that decompose the emotion-cause pair extraction process into two steps. The first step involves predicting the emotions of the utterances in the conversation. In the next step, we utilize these emotion labels to guide cause extraction. Approach 1 involves instruction-tuning two separate Llama 2 models for emotion and cause prediction, while Approach 2 leverages the in-context learning (ICL) capabilities (Dong et al., 2023) of the GPT-3.5 model. Additionally, we introduce an efficient technique using the GPT-4V model to extract conversation-level descriptions from video modality.
@@ 2.1 Task definition
2.2 Related Work
The detailed Related Work section can be found in Appendix  A.
2.3 Dataset
@@
We use the Emotion-Cause-in-Friends (ECF) dataset provided by Wang et al. (2023a), which is summarized in Table 1. This dataset contains 13,509 multimodal utterances that occur in the American sitcom Friends with 9272 emotion-cause pairs. Each utterance consists of the text, video, and audio.
@@
3 Methodology
3.1 Overview
We treat the task at hand as a two-step process. In the first step, we predict the emotion of each utterance in all N conversations. Here, the context $C_{j}$ for utterance $u_{j}$ of conversation $D_{i}$ is the entire conversation itself. Given $E$ target emotion labels and $\hat{y_{j}^{e}}$ as the predicted emotion label, the problem can be formulated as (where $\theta$ denotes the parameters):
$\hat{y_{j}^{e}}=\arg\max_{e}\mathcal{P}(y^{e}|u_{j},C_{j},\theta)$ (1)
In the second step, given these emotion labels, we predict the causes of each utterance that has an emotion other than neutral. The causes will be a subset of all utterances in the conversation $D_{i}$. Let the learned function be $f:U\rightarrow 2^{U}$, where $U$ is the set of all utterances in the given conversation. It predicts the subset $\hat{y_{j}^{c}}$ of cause of emotion utterance $u_{j}$ where $\hat{y_{j}^{e}}\neq$ neutral as:
3.2 Approach 1: Fine-tuned Llama-2
In our first approach, we perform instruction fine-tuning of the Llama 2 Large Language Model, an open-source model developed by GenAI, Meta (Touvron et al., 2023). From the three variants with 7, 13, and 70 billion parameters, we use the 13 billion parameter model due to resource constraints, albeit the performance of this model achieves state-of-the-art results on various downstream NLP tasks compared to other models of similar sizes (Touvron et al., 2023). In addition, we use the Llama 2-chat version of the model [^1] https://huggingface.co/meta-llama/Llama-2-13b-chat-hf, which is optimized for dialogue use cases as it aligns with our task. In our approach, we use Llama2 API [^2] https://www.llama2.ai/ for prompt engineering. Through zero-shot prompting, we select optimal prompts for emotion identification and cause prediction. We observed that treating these two tasks separately resulted in better model output. This approach involves first identifying the emotions of all utterances in the conversation. We then add these emotion labels to the conversation and prompt the model to predict the causes for each emotion utterance. Consequently, we perform supervised fine-tuning of two separate Llama 2 models for these tasks. Although this increases the inference time, the significant performance gains outweigh the introduced latency. We treat both tasks as conditional generation, where the model generated the emotion label in the first case and the cause list in the second case, given the prompt. Detailed explanations of these approaches are provided in the following sections. The fine-tuning procedure is shown in Fig.3.
Figure 3: Pipeline for fine-tuning Llama (Approach 1)
3.2.1 Emotion recognition
To perform emotion recognition, we create a dataset where each sample includes an utterance $u_{j}$ from one of the $N$ conversations $D$ for which the LLM needs to output the emotion label. We incorporate the entire conversation $D_{i}$ along with speaker information as context in our prompt. This contextual information enhances the model’s understanding of the flow of emotions within the conversation, as demonstrated by our ablation studies in Section 5. The instruction $I_{j}^{e}$, which gave the best results, is given in Appendix  E.1 along with detailed prompt examples. The prompt consists of the instruction $I_{j}^{e}$ and the context $C_{j}$ for utterance $u_{j}$:
$Prompt_{j}=(C_{j},I_{j}^{e})$ (3)
$\hat{y_{j}^{e}}=\textbf{Llama}_{e}(Prompt_{j},\theta)$ (4)
3.2.2 Cause prediction
To prepare the dataset for cause prediction, we incorporate the emotion labels obtained for each utterance. The conversation context now includes the emotion labels for each utterance $u_{j}$ excluding those with a predicted emotion label $\hat{y_{j}^{e}}$ of neutral. This approach enhances the model’s ability to analyze causal dependencies and identify which utterances may have contributed to a specific emotion. The output for cause prediction is a list of cause utterance IDs. The instruction is provided in Appendix  E.1. The modified prompt for this step consists of this instruction $I_{j}^{c}$ along with the conversational context with emotion labels $C_{j}^{e}$:
$Prompt_{j}=(C_{j}^{e},I_{j}^{c})$ (5)
$\hat{y_{j}^{c}}=\textbf{Llama}_{c}(Prompt_{j},\theta)$ (6)
3.2.3 Adding video captions
3.3 Approach 2: In-Context-Learning GPT
Figure 4: Pipeline of In-Context-Learning GPT Method (Approach 2)
3.3.1 Video Captioning
Figure 5: Video Captioning Pipeline
3.3.2 Emotion Recognition
GPT tends to be uncontrollable when performing zero-shot recognition of emotions in conversations Qin et al. (2023) outputting emotions that are not a valid category of labels. To guide and control the process, we leverage in-context learning (ICL) by retrieving a conversation from the training set whose emotions are already annotated. The emotions in these conversations are explained by GPT-3.5 (Appendix Fig. 17). This retrieved conversation and its explanation serve as a demonstration for GPT to learn from, enabling it to recognize emotions in conversations more accurately. In addition, the prompt template includes the video caption as part of the input, as shown in Appendix Fig. 18.
To ensure effective ICL, it is important to provide general and descriptive examples that aid in solving the current task. In our approach, we sampled conversations from the training set containing all emotion categories. These conversations were stored as text-embedding-ada-002 embeddings Neelakantan et al. (2022) in a vector database. At test time, we compute the embedding for a conversation and retrieve the closest matching embedding from the database based on Euclidean distance. The retrieved embedding aids ICL in improving emotion understanding and recognition.
3.3.3 Cause Prediction
Following the prediction of emotions, we predict the causes for each emotional utterance within a context window around that utterance. The bounds of the context window are given in Table 2. The bounds were informed by the distribution of the majority of relative positions of causes in the training set (Figure 2).
Position Previous Next
Beginning 0 2
End 5 0
Middle 5 2
Table 2: Context Window Bounds in each Direction
For predicting the causes of an utterance with emotion $e$ within a given context window $c$, we retrieve context windows containing utterances with the same emotion $e$ that exhibit semantic similarity to $c$. This retrieval is accomplished through the Euclidean distance comparison of text-embedding-ada-002 embeddings derived from the training data. The retrieved conversation’s causes are explained by GPT-3.5 (Appendix Fig. 19). Learning from the explained retrieved-context windows, cause prediction on $c$ can be performed by GPT-3.5. Video captions are also included in the prompt (Appendix Fig. 20), since the local window may have lost some broader context.
3.4 Post-Processing
In both our approaches, after getting the causes, we perform a post-processing step where we add the emotional utterance as its own cause which we call self-causes. This gives significant performance boosts as a majority of the causes are self-causes as pointed out in Appendix  2.3.
4 Experimental setup
Training details
For approach 1, the data is split into train, test, and validation sets in the ratio 8:1:1. We use peft library  [^4] https://huggingface.co/docs/peft/en/index for Parameter-Efficient Fine-Tuning. Due to memory constraints, we fine-tune a 4-bit quantized Llama-2 model using bitsandbytes library  [^5] https://github.com/TimDettmers/bitsandbytes. We report the details of the implementation for both approaches in Appendix  B.
Evaluation metrics
5 Results and Discussion
Main results
System w-avg F1 F1
1. Samsung Research China-Beijing 0.3774 0.3870
2. NUS-Emo 0.3460 0.3517
3. SZTU-MIPS 0.3435 0.3434
4. GPT-ICL (Ours) 0.2758 0.2816
5. MotoMoto 0.2584 0.2595
6. Fine-tuned Llama (Ours) 0.2558 0.2630
Table 3: Leaderboard Results on Evaluation Data
Ablation study
We conduct extensive ablation studies to measure the importance of the techniques we employ summarized in Table  4. For these experiments, we use a subset of our test set containing 528 utterances. It can be seen that the performance of zero-shot Llama as well as GPT is the lowest. Instruction-tuning and ICL clearly improve the performance on the task, showcasing the significance of making LLMs context-aware when tackling downstream tasks. Adding self-causes improves performance in both zero-shot and context-aware cases highlighting their importance. The incorporation of video captions leads to poorer results in context-learning. The detailed table is in Appendix  C.
Approach F1 w-avg F1
Zero-shot Llama
      - w/o self-causes 0.117 0.116
      - w/ self-causes 0.222 0.215
Instruction-tuned Llama
      - w/o self-causes 0.325 0.318
      - w/ self-causes 0.364 0.352
Zero-shot GPT
      - w/o self-causes 0.100 0.097
      - w/ self-causes 0.189 0.184
In-context-learning GPT
      - w/o self-causes w/o video 0.286 0.296
      - w/o self-causes w/ video 0.235 0.241
      -w/ self-causes w/o video 0.336 0.342
      -w/ self-causes w/ video 0.329 0.334
Table 4: Results on Validation Set.
Limitations
6 Conclusion
Appendix A Related Work
Text-based ECA
Multimodal ECA
Large Language Models
Appendix B Implementation details
B.1 Training details for Llama
@@
B.2 Details for in-context learning GPT
@@
In-context-learning GPT w/o video captions w/o self-causes 0.259 0.319 0.286 0.283 0.319 0.296
In-context-learning GPT w/o video captions w/ self-causes 0.270 0.445 0.336 0.287 0.445 0.342
@@
Appendix C Detailed results
@@
Appendix D Error Analysis
@@
Appendix E Prompt details
E.1 Fine-tuned Llama 2